# Agentic Multi-Persona Framework for Evidence-Aware Fake News Detection

### v1

@@ 5.3. Ablation Study
(i) Impact of Persuasion Analysis (PKM): PKM alone achieved relatively weak performance ($Acc$: 73.13 on PolitiFact, 67.90 on GossipCop), confirming that persuasion cues by themselves cannot reliably classify misinformation. However, ablating PKM from the full system (i.e., w/o PKM) causes notable drops (e.g., –20.66% $Acc$ on PolitiFact, –37.60% on GossipCop), highlighting its value in resolving uncertain cases where factual evidence is ambiguous. Thus, PKM is complementary: it does not excel standalone but adds critical discriminative power when integrated.
(ii) Impact of Evidence Base: Evidence-only models ($Acc$: 74.26 on PolitiFact, 79.67 on GossipCop) establish factual signals as a solid baseline but fall short compared to more integrated approaches. The large gains when evidence was combined within AMPEND (e.g., +20.61% $F1$ on PolitiFact) show that credibility scoring, temporal filtering, and multimodal grounding are indispensable for anchoring agentic reasoning to verifiable facts.
(iii) Impact of Agentic Reasoning: The standalone agentic framework outperformed PKM and Evidence individually ($Acc$ 79.92 PolitiFact, 83.14 GossipCop), demonstrating its ability to structure reasoning and probe claims. However, without external modules, it still underperformed compared to the full system. This indicates that while reasoning agents can organize analysis, they need persuasion and evidence signals for robustness.
(iv) ASPEND vs. AMPEND: A single-persona reasoning model, ASPEND, achieved strong performance ($Acc$ 84.13 PolitiFact, 88.80 GossipCop), but lagged behind the multi-persona AMPEND ($Acc$: 88.54 PolitiFact, 92.68 GossipCop). This confirms that multi-persona reasoning captures diverse perspectives (legal, journalistic, scientific, supervisory) that a single persona cannot, leading to more reliable and nuanced verdicts.
(v) AMPEND vs. AMPEND-LS: The final hybrid system AMPEND-LS, integrating the complementary SLM, yields further improvements ($Acc$: 92.18 PolitiFact, 93.43 GossipCop). The gains, though modest over AMPEND, reflect significant stability: the SLM acts as a lightweight decision layer that mitigates LLM uncertainty and hallucinations, ensuring balanced precision–recall and reducing overfitting. This synergy between LLM reasoning and SLM classification is crucial for both scalability and reliability.
@@
Figure 2. Performance across multi-rounds Q&A on PoltiFact.
5.4. Sensitivity of the Number of Rounds for Efficient Multi-Persona Reasoning
Figure 2 shows that AMPEND-LS exhibits stable yet incremental performance gains across multiple rounds of multi-persona Q&A on PolitiFact: $Acc$ improves from 88.54 at round-1 to 89.30 at round-4 (+0.76%), $F1$ from 87.34 to 88.54 (+1.20%), $Rec$ from 86.32 to 87.82 (+1.50%), while $Pre$ remains consistently high ($\sim$90%) with minimal fluctuations. These results highlight that each additional round contributes positively by refining evidence and balancing precision–recall trade-offs; however, the marginal improvements diminish after round 3. Since multi-agent Q&A cycles rely on costly API calls to LLMs, continuing indefinitely is neither practical nor efficient. Thus, a sweet spot can be identified around 2–3 rounds, where significant gains in accuracy and $F1$ are already realized, while avoiding unnecessary computational and monetary overhead. This balance between performance improvements and resource efficiency underscores the practicality of limiting reasoning depth while preserving robustness in real-world deployments.
5.5. Information-Theoretic Evaluation of LLM-Justification
To assess the effectiveness of LLM-generated justifications, we evaluate their impact across three complementary dimensions: $REV$ (rev_metric), $LAS$ (las_metric), and the utility performance metrics $\mathcal{U}_{i}$, where $i\in\{\text{$Acc$},\text{$F1$},\text{$Pre$},\text{$Rec$}\}$, as reported in Table 4. Across all datasets, our model exhibits consistently positive $REV$ and $LAS$ scores, demonstrating that LLM justifications contribute additional semantic information (higher $REV$) while maintaining faithful alignment with model predictions (positive $LAS$). More importantly, the utility metrics show notable performance gains, indicating that integrating LLM rationales into the SLM finetuning effectively enhances both predictive accuracy and explanatory reliability.
Among all datasets, PolitiFact achieves the largest improvements, suggesting that factual contexts benefit more from explicit justification signals, potentially due to the causal alignment between evidence and claims. GossipCop and MMCoVaR also show steady improvements, highlighting the AMPEND-LS adaptability across different content domains and rationale granularities. These results underscore that faithful and informative LLM justifications can meaningfully improve downstream task performance, bridging the gap between explainability and utility in aligned SLM models.
Table 4. LLM justification analysis
@@ 5.5. Information-Theoretic Evaluation of LLM-Justification
6. Limitations and Future Work
While AMPEND-LS achieves strong performance and interpretability across diverse benchmarks, several limitations highlight opportunities for future extensions. The framework’s reliance on external APIs (e.g., web and reverse image search) may introduce latency, coverage gaps, or dependency risks; future work can investigate retrieval-agnostic or locally cached evidence pipelines to ensure robustness. The multi-persona reasoning mechanism, though improving accuracy and transparency, requires repeated LLM interactions that increase computational and monetary cost. Designing adaptive cost-aware policies to dynamically control reasoning depth offers a promising direction for scalable real-time deployment. Moreover, our experiments are limited to three English-centric datasets (PolitiFact, GossipCop, MMCoVaR). Extending evaluation to multilingual, low-resource, and culturally diverse misinformation scenarios is a critical step toward broader applicability. Current credibility scoring draws on third-party databases such as MBFC, which may not capture evolving global media ecosystems; adaptive credibility estimation that updates with new outlets can strengthen long-term resilience. Finally, while persuasion detection and evidence fusion enhance robustness, adversarial manipulations such as AI-generated deepfakes or coordinated campaigns remain an open frontier. Incorporating adversarial training, generative counterfactuals, and stronger multimodal alignment will further improve resilience.
@@ 6. Limitations and Future Work
By acknowledging these challenges as natural extensions, we emphasize that AMPEND-LS provides a solid and adaptable foundation. Addressing the above directions will not only enhance scalability and inclusivity but also strengthen the system’s role as a deployable, trustworthy solution for combating misinformation in dynamic, real-world environments.
7. Conclusion
@@
This work introduced AMPEND-LS, an agentic multi-persona framework for evidence-aware fake news detection that unifies multimodal evidence retrieval, structured reasoning, and hybrid classification. By combining LLM-driven multi-perspective analysis with lightweight SLM adaptation, the framework achieves state-of-the-art performance while offering transparent and scalable decision-making. Extensive experiments across political, entertainment, and health misinformation domains confirm its robustness and interpretability. Looking ahead, the modular design provides a strong foundation for advancing toward multilingual, low-resource, and adversarially robust misinformation defense, thereby paving the way for trustworthy deployment in real-world, high-stakes information ecosystems.
Appendix A Configuration Details
Table 5 presents the hyperparameters of AMPEND-LS. The top document counts ($K_{i}$ or $K^{\prime}_{i}$) and the number of KG paths are set empirically to maintain relevance while reducing noise. Domain credibility thresholds align with MBFC category labels, and the temporal filter $\alpha_{t}$ excludes documents published within roughly a year. The number of persona Q&A steps is tuned using the PolitiFact dataset, while the LLM temperature is adjusted to balance creative generation with focused inference.
Table 5. Model configuration and hyperparameters
Attribute Value Top document Counts $K_{1}=K_{2}=15,K^{\prime}_{1}=K^{\prime}_{2}=10$ Domain Credibility Score $\alpha_{h}=0.9,\alpha_{m}=0.7,\alpha_{u}=0.5,\alpha_{l}=0.3$ Temporal Evidence Filtering $\alpha_{t}=0.2$ # of KG Triplet 10 # of Rounds ($\tau$) 4 Temperature Claim generation 0.30 KG generation 0.20 Persuasion generation 0.30 Question generation 0.70 Answer generation 0.40 Memory Update 0.50 LLM Classifier 0.25
Table 6. Qualitative Analysis: Correctly Classified Cases
Image Headline Article Content LLM Justification Ground Truth [Uncaptioned image] Air India Flight 171 Air India Flight 171 was a scheduled passenger flight from Ahmedabad Airport in India to London Gatwick Airport in the United Kingdom that crashed 32 seconds after takeoff at 13:39 IST (08:09 UTC) on 12 June 2025. All 12 crew members and 229 of the 230 passengers aboard died. On the ground, 19 people were killed and 67 others were seriously injured…. Verdict: REAL; Confidence: 95 Justification: The claim regarding the crash of Air India Flight 171 on June 12, 2025, is confirmed by multiple credible sources, including a preliminary report from India’s Aircraft Accident Investigation Bureau (AAIB) and coverage from medium-credibility sources like Wikipedia (wikipedia.org) and Aerotime (aerotime.aero) . These sources provide specific details about the crash, including the loss of thrust due to fuel control switches moving to CUTOFF, corroborating the tragic loss of life and the ongoing investigation. REAL [Uncaptioned image] Cristiano Ronaldo Joining Lionel Messi at Inter Miami for Club World Cup? In a surprising turn of events, football legends Cristiano Ronaldo and Lionel Messi are reportedly set to team up at Inter Miami for the upcoming Club World Cup. Sources close to the club have hinted at a groundbreaking partnership between the two superstars, aiming to elevate the team’s performance on the global stage… Verdict: FAKE; Confidence: 95
Justification: Cristiano Ronaldo remains under contract with Al Nassr, and there is no credible evidence supporting his transfer to Inter Miami. Reports suggesting this potential move are speculative and lack official confirmation from reputable sources. Messi’s confirmed presence at Inter Miami is the only verified aspect of the claim. Key sources include: (ESPN) and (NPR).
FAKE
Table 7. Qualitative Analysis: Misclassified Cases
Image Headline Article Content LLM Justification Ground Truth [Uncaptioned image] Ohio student suspended for staying in class during National Walkout Day An Ohio high school student said he tried to stay apolitical during the National Walkout Day over gun violence but was suspended for his choice to remain in a classroom instead of joining protests or the alternative, going to study hall… Verdict: REAL; Confidence: 95 Justification: Multiple high-credibility sources, including Snopes (snopes.com) and Politifact (politifact.com), confirm that Jacob Shoemaker was indeed suspended for remaining in class during National Walkout Day. The reports clarify that while he did not participate in the protest, he was required to be supervised, which led to his suspension. This aligns with school policies on student conduct during organized events. FAKE [Uncaptioned image] Ben Affleck & Lindsay Shookus: Planning a Wedding Even Though He’s Still Married?! It’s been almost a year since the world first learned that Ben Affleck and Lindsey Shookus are dating, and these days, the couple is reportedly ready to take the next step together.Unfortunately, there are some sizable barriers blocking their path. Verdict: UNCERTAIN; Confidence: 70 Justification: The claim about Ben Affleck and Lindsay Shookus planning a wedding is based on reports from various sources, but there is no definitive confirmation from high-credibility outlets. While some sources suggest they are in a serious relationship, the ongoing legal marriage of Affleck to Jennifer Garner complicates the situation. Additionally, there are conflicting reports regarding Affleck’s readiness to propose, such as Us Magazine (usmagazine.com) and E! Online (eonline.com), leading to uncertainty about the claim’s validity. FAKE [Uncaptioned image] Peter Strzok, FBI Agent in Texting Scandal, Will Testify Before Congress Peter Strzok, the FBI agent who sent anti-Trump messages to a colleague while leading the investigation of Hillary Clinton’s emails and Russia’s involvement in the Trump campaign, has said he is willing to testify before the House Judiciary Committee, or any other committee that wants to speak with him… Verdict: REAL; Confidence: 90 Justification: The article accurately describes Peter Strzok’s willingness to testify before Congress regarding his involvement in the investigations into Hillary Clinton’s emails and Russian interference in the Trump campaign. This is supported by multiple high-credibility sources, including reports from CBC (cbc.ca), PBS (pbs.org), and Vox (vox.com), confirming Strzok’s testimony and the context of his anti-Trump messages. The details align with the findings of the DOJ inspector general’s report, which found no evidence that Strzok’s personal views influenced specific investigative decisions. FAKE
Appendix B Qualitative Analysis
To complement the quantitative evaluation, we conducted a qualitative study to analyze how AMPEND-LS performs on individual news instances. Importantly, the examples considered here include cases drawn from outside the benchmark datasets, demonstrating the framework’s ability to generalize to real-world scenarios. The analysis highlights not only correctly classified instances but also misclassifications, thereby providing a balanced perspective on the system’s interpretability and limitations.
B.1. Correctly Classified Cases
Correct classifications occur when the model’s predictions align with the ground truth through effective evidence retrieval and structured reasoning.
Verifiable factual reporting supported by authoritative sources: In the case of the Air India Flight 171 crash (June 12, 2025), AMPEND-LS correctly predicted REAL, corroborating the claim with multiple reliable sources, including India’s Aircraft Accident Investigation Bureau and aviation-focused outlets. The LLM justification emphasized consistent technical details (e.g., fuel control switches moving to CUTOFF) and confirmed casualties, reflecting disciplined evidence selection and precise grounding in domain-specific facts.
Speculative or sensational claims lacking credible confirmation: For the Cristiano Ronaldo transfer rumor to Inter Miami, the model correctly classified the article as FAKE, citing Ronaldo’s ongoing contract with Al Nassr and the absence of credible confirmation. The justification explicitly contrasted Messi’s verified presence at Inter Miami with the unsupported rumor of Ronaldo joining, illustrating the framework’s ability to detect fabricated narratives despite their surface plausibility.
These cases demonstrate that the LLM-generated justifications are not only concise and auditable but also grounded in verifiable evidence, reinforcing transparency and user trust.
B.2. Misclassified Cases
Despite strong overall performance, some cases revealed inherent challenges where the model’s reasoning diverged from dataset labels.
Context-sensitive claims: The Ohio student suspension during National Walkout Day was predicted as REAL with high confidence because outlets like Snopes and PolitiFact confirmed the suspension event. However, the dataset labeled it FAKE, since the suspension was linked to procedural supervision rules rather than political protest. This case highlights the difficulty of handling claims where factual correctness diverges from contextual intent.
Conflicting but credible evidence: In the Ben Affleck–Lindsay Shookus wedding rumor, AMPEND-LS output UNCERTAIN, reflecting the mixed evidence across entertainment outlets. While the dataset labeled it FAKE, the model’s cautious stance demonstrates its ability to avoid overcommitment in ambiguous settings, though benchmark scoring penalizes such conservative predictions.
Framing effects versus factual content: For Peter Strzok’s congressional testimony, the model predicted REAL, supported by reputable outlets such as CBC, PBS, and Vox. Yet the dataset labeled it FAKE, likely due to framing or rhetorical exaggeration in the original claim. Here, AMPEND-LS prioritized factual accuracy over political tone, revealing tension between annotation policy and evidence-based reasoning.
Discussion: Taken together, the qualitative study illustrates that AMPEND-LS’s strengths lie in its ability to validate factual claims through evidence-grounded reasoning and to detect fabricated or speculative narratives. Misclassifications, however, reveal limitations in handling (i) context-sensitive misinformation where truth and intent diverge, (ii) ambiguity arising from credible yet conflicting sources, and (iii) discrepancies between dataset labeling criteria and factual verification. Crucially, the LLM justifications in both correct and incorrect cases remain coherent, auditable, and evidence-aware, underscoring that errors stem from reasonable interpretations rather than hallucination or arbitrary misjudgment.
Overall, the qualitative analysis confirms that AMPEND-LS not only improves predictive performance but also provides transparent reasoning, making its decisions interpretable to both users and researchers. These insights further motivate future refinements, such as context-aware reasoning, explicit uncertainty modeling, and dataset designs that better capture rhetorical framing alongside factual accuracy.
Table 8. Cross-domain analysis
Training Data Test Data Method AccAcc $F1$ PrePre RecRec PolitiFact GossipCop MGCA 49.79 49.65 49.67 49.65 BREAK 55.43 55.08 59.87 51.00 AMPEND-LS 85.33 85.43 85.53 85.33 MMCoVaR BREAK 60.48 60.46 60.13 60.79 MGCA 55.88 52.52 52.58 52.46 AMPEND-LS 76.25 75.89 81.83 70.68 GossipCop PolitiFact MGCA 43.41 43.36 41.89 44.94 BREAK 56.97 56.92 58.35 55.55 AMPEND-LS 82.78 82.65 84.67 80.73 MMCoVaR MGCA 50.18 49.66 49.56 49.76 BREAK 61.12 57.29 58.27 56.34 AMPEND-LS 80.58 80.21 79.85 80.58 MMCoVaR PolitiFact MGCA 41.94 43.92 43.81 44.04 BREAK 71.27 68.01 69.79 66.32 AMPEND-LS 83.09 84.46 85.87 83.09 GossipCop BREAK 53.40 53.45 53.51 53.40 MGCA 53.53 53.48 53.74 53.22 AMPEND-LS 74.21 73.95 74.22 73.68 PolitiFact + GossipCop MMCoVaR MGCA 62.37 60.02 61.49 58.62 BREAK 59.45 54.71 58.77 51.17 AMPEND-LS 81.57 82.85 84.17 81.57 PolitiFact + MMCoVaR GossipCop BREAK 61.60 61.63 61.66 61.60 MGCA 64.28 64.58 63.58 65.61 AMPEND-LS 71.43 71.06 71.03 71.10 GossipCop + MMCoVaR PolitiFact BREAK 68.45 65.56 68.43 62.92 MGCA 70.56 70.60 77.32 64.95 AMPEND-LS 79.41 79.01 84.40 74.26 - PolitiFact AMPEND 88.54 88.25 90.26 86.32 GossipCop 92.68 92.57 93.04 92.10 MMCoVaR 91.27 89.19 92.11 86.45
Appendix C Cross-domain Analysis
To rigorously assess the domain generalization capability of our framework, we evaluate the framework across multiple datasets that capture diverse misinformation domains. Table 8 reports comparisons against two strong baselines: MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) and BREAK (yin2025break). MGCA represents state-of-the-art multimodal approaches, aligning textual and visual signals via cross-modal contrastive learning, while BREAK exemplifies transformer-based domain adaptation through structure- and attention-aware mechanisms. These baselines were selected not only for their competitiveness in prior literature but also because they consistently performed well in our experiments, often ranking as the second-best methods across settings. This makes them appropriate and challenging comparators for evaluating the robustness of AMPEND-LS under distribution shift. The evaluation spans three heterogeneous datasets: PolitiFact (political claims), GossipCop (entertainment news), and MMCoVaR (COVID-19 health misinformation), which together reflect the topical diversity, stylistic variation, and domain heterogeneity characteristic of real-world misinformation.
[…]
(iv) AMPEND vs. AMPEND-LS: Table 8 also provides a comparative view of domain generalizability between AMPEND and AMPEND-LS. AMPEND, which directly applied the full multi-persona reasoning pipeline without distillation, exhibited stronger cross-domain generalization than AMPEND-LS. This was expected, as AMPEND leveraged rich LLM-based reasoning and iterative evidence aggregation without compression, allowing it to adapt more effectively to unseen distributions. However, when training and testing domains coincided, AMPEND-LS achieved higher performance than AMPEND, since the SLM distillation captured and reinforced domain-specific patterns while retaining evidence-grounded signals. Importantly, despite a slight degradation relative to AMPEND in true cross-domain transfer, AMPEND-LS still demonstrated substantially higher generalizability than competitive baselines such as MGCA and BREAK.
Overall, these results establish that AMPEND-LS achieves domain-agnostic robustness, balanced precision–recall, and scalable efficiency, outperforming competitive baselines under both single-source transfer and multi-source training settings. At the same time, AMPEND serves as a reasoning backbone for maximal cross-domain generalization, while AMPEND-LS extends this capability into a practical, efficient framework suitable for real-world misinformation detection.

### v2

@@ V-C Ablation Study
MMCoVaR PKM 71.94 63.10 76.21 72.20
Evidence 80.04 74.10 78.12 80.02
Agentic 81.20 76.01 78.15 81.10
w/o PKM 85.97 85.02 86.20 86.01
ASPEND 86.28 85.12 85.01 86.20
AMPEND 91.27 89.19 92.11 86.45
AMPEND-LS 96.46 95.56 96.30 94.96
@@
Figure 2: Performance across multi-rounds Q&A on PoltiFact.
V-D Sensitivity to Reasoning Depth in Multi-Persona Framework
Fig. 2 illustrates the impact of increasing the number of multi-persona Q&A rounds on the PolitiFact dataset. AMPEND-LS exhibits stable and incremental performance improvements as reasoning depth increases. Specifically, $Acc$ improves from 88.54 at round 1 to 89.30 at round 4 (+0.76 percentage points), $F1$ increases from 87.34 to 88.54 (+1.20 pp), and $Rec$ rises from 86.32 to 87.82 (+1.50 pp). Meanwhile, $Pre$ remains consistently high (approximately 90%) with only minor fluctuations across rounds. These results indicate that additional reasoning cycles help refine evidence aggregation and improve the precision–recall balance. However, the marginal performance gains diminish beyond the third round. Since each reasoning cycle incurs additional LLM calls, increasing the number of rounds directly impacts latency and computational cost. From a systems perspective, 2–3 rounds provide a favorable trade-off, capturing most of the performance gains while avoiding unnecessary overhead. This controlled reasoning-depth analysis further supports the deployment design of AMPEND-LS, where adaptive round selection can dynamically balance accuracy, robustness, and computational efficiency in real-time misinformation monitoring scenarios. Although only PolitiFact is shown here for clarity, we conducted analogous experiments on GossipCop and MMCoVaR and observed consistent trends, confirming that the reasoning-depth trade-off generalizes across datasets.
V-E Information-Theoretic Evaluation of LLM-Justification
To assess the effectiveness of LLM-generated justifications, we evaluate their contribution along three complementary dimensions: $REV$ [3], $LAS$ [10], and the utility performance gains $\mathcal{U}_{i}$, where $i{\in}\{\text{$Acc$},\text{$F1$},\text{$Pre$},\text{$Rec$}\}$, as reported in Table III. Across all datasets, AMPEND-LS achieves consistently positive $REV$ and $LAS$ scores, indicating that the generated justifications provide additional label-relevant information (higher $REV$) while maintaining faithful alignment with model predictions (positive $LAS$). Furthermore, all utility metrics are positive, demonstrating that incorporating LLM-generated rationales into SLM fine-tuning yields measurable improvements in predictive performance. Among the three datasets, PolitiFact exhibits the largest gains in both information-theoretic measures and downstream utility (e.g., $\mathcal{U}_{F1}=11.72$, $\mathcal{U}_{Rec}=13.62$), suggesting that justification signals are particularly beneficial in fact-centric political contexts. GossipCop and MMCoVaR also show consistent, though comparatively moderate, improvements, confirming that the benefits of justification integration generalize across entertainment and health misinformation domains. Overall, these findings show that faithful and informative LLM justifications contribute meaningful task-relevant signals that simultaneously improve downstream performance and interpretability, effectively bridging explainability and utility in aligned SLM models.
TABLE III: LLM justification analysis
@@ V-E Information-Theoretic Evaluation of LLM-Justification
Appendix A outlines the configuration details of AMPEND-LS, Appendix B presents the qualitative analysis, and Appendix C provides the cross-domain analysis.
VI Limitations and Future Work
While AMPEND-LS achieves strong performance and interpretability across diverse benchmarks, several limitations point to promising directions for future research. The framework’s reliance on external APIs (e.g., web and reverse image search) may introduce latency, coverage gaps, or dependency risks. Future work may explore retrieval-agnostic or locally cached evidence pipelines to improve robustness and deployment reliability.
The multi-persona reasoning mechanism, although enhancing accuracy and transparency, requires repeated LLM interactions that increase computational and monetary cost. Designing adaptive, cost-aware policies to dynamically control reasoning depth presents a promising avenue for scalable real-time deployment.
Furthermore, our experiments are limited to three English-centric datasets (PolitiFact, GossipCop, and MMCoVaR). Extending evaluation to multilingual, low-resource, and culturally diverse misinformation scenarios is essential for broader applicability. Current credibility scoring relies on third-party databases such as MBFC, which may not fully capture the dynamics of evolving global media ecosystems; adaptive credibility estimation that continuously incorporates emerging sources could strengthen long-term resilience.
Finally, although persuasion detection and multimodal evidence fusion enhance robustness, adversarial manipulations—such as AI-generated deepfakes and coordinated misinformation campaigns—remain open challenges. Incorporating adversarial training, generative counterfactual reasoning, and stronger multimodal alignment mechanisms may further improve system resilience.
@@ VI Limitations and Future Work
By acknowledging these limitations as natural extensions rather than shortcomings, we emphasize that AMPEND-LS provides a solid and adaptable foundation. Addressing these directions will enhance scalability, inclusivity, and robustness, strengthening the system’s role as a deployable and trustworthy solution for combating misinformation in dynamic real-world environments.
VII Conclusion
@@
This work introduced AMPEND-LS, an agentic multi-persona framework for evidence-aware fake news detection that integrates multimodal evidence retrieval, structured reasoning, and hybrid LLM–SLM classification within a unified architecture. By combining multi-perspective, role-aware analysis with lightweight SLM adaptation, the framework achieves state-of-the-art performance while maintaining transparency, robustness, and scalability. Extensive experiments across political, entertainment, and health misinformation domains demonstrate consistent improvements in predictive accuracy and interpretability. The modular design further provides a strong foundation for extension to multilingual, low-resource, and adversarially robust settings, paving the way for trustworthy deployment in real-world, high-stakes information ecosystems.
Supplementary Appendix
Appendices A, B, and C are provided in https://github.com/csksuraj17/AMPEND-LS_Supp